# from Monocular RGB-D Video

### v1

@@ 1 Introduction
The challenge of reconstructing the dense 3D geometry, appearance, and motion of dynamic scenes from videos has persisted for decades, offering a wide range of applications in virtual reality and augmented reality. Due to the inherent ambiguities of the problem, traditional dense non-rigid reconstruction approaches often assume strong priors on the object’s shape and motion. Model-based approaches [75, 1, 26, 11, 65, 25] require 3D parametric models [2, 36, 42, 6, 28, 84] of specific object categories such as humans [2, 36], faces [6, 28], or animals [84]. While these methods achieve promising reconstruction results, they fall short of generalizing to arbitrary object categories. In this paper, we focus on model-agnostic approaches that could generalize to generic non-rigid objects. This task is often approached by jointly fusing a canonical shape and a per-frame deformation field from an RGB-D video sequence [41, 24, 59, 60, 8, 7]. Due to the inherent constraints of physical sensors and the use of traditional voxel-grid representations, they often struggle to produce high-fidelity surface details and can only reconstruct observed regions.
Recent attention has shifted towards leveraging neural representations for scene reconstruction. Neural Radiance Fields (NeRF) [39] represent scene density and color in the weights of a neural network. In combination with volume rendering, NeRF achieves unprecedented performance on novel view synthesis. Many follow-up works adapt this idea into surface reconstruction [79, 72, 4, 71], SLAM [61, 82, 27], dynamic scene reconstruction [48, 44, 45, 9] and more. Notably, NDR [9] proposes the first RGB-D neural implicit dynamic surface reconstruction method. Thanks to the expressiveness of neural representations, they achieve accurate and smooth surface reconstruction. However, real-world captures often contain large unobserved regions. The inherent smoothness of neural representations, and the difficulty to incorporate domain-specific priors, causes current methods to fall short of achieving realistic completion.
@@
Our key contribution is to demonstrate the capability to learn metrically accurate geometry and deformations of dynamic objects from casually captured RGB-D videos while achieving realistic completion in unobserved regions with diffusion priors. We propose a temporal view-dependent SDS strategy to achieve photo-realistic completion while learning an accurate deformation field. A canonical space regularization strategy is used to avoid the trivial solution of surface completion being explained via deformations. Experimental results across various datasets show that MorpheuS achieves competitive reconstruction accuracy and realistic completion.
@@ 2 Related Work
Dense Non-rigid Reconstruction. Traditional dense dynamic scene reconstruction methods often require strong prior knowledge about the object shape and motion, fitting 3D prior parametric models [36, 51, 84, 28, 75, 1, 26, 11, 65, 25] to video sequences of a specific object category, and thus cannot work with arbitrary unseen objects. Early model-agnostic methods either require accurate dense correspondences [19] or a template shape of the object to be reconstructed first before tracking the dynamic motion [62, 83, 80, 5]. The seminal work DynamicFusion [41] was the first direct and template-free method to achieve real-time simultaneous tracking and reconstruction of generic non-rigid objects via joint optimization of a canonical shape and a per-frame deformation field. Follow-up methods improve the robustness of motion tracking and reconstruction quality by incorporating photo-metric constraints [20], classical  [24] or learned correspondences [8, 7], reformulating the optimization as direct dense SDF alignment [59, 60]. Although these methods have shown promising reconstruction results, they fall short of completing the unobserved regions due to their inherent traditional scene representation and lack of generative priors.
@@
Neural Rendering for Dynamic Scenes. The success of NeRF [39] in novel view synthesis has inspired many follow-up works [48, 44, 29, 45, 17, 66, 68, 15, 9, 18, 31] in dynamic scene reconstruction. These approaches can be broadly classified into two paradigms: a) Directly modeling the scene geometry and motion as a unified 4D space-time field in the world (observation) space [29, 17, 10, 16, 31, 54, 23], where motion priors, such as scene flow or trajectory can be involved as an additional regularization [29, 68, 31]. b) Learning a canonical shape by warping the shape from observation space to canonical space via a separate deformation field [48, 44, 45, 66, 15, 9, 18, 76, 14]. Among these works, NDR [9] is particularly relevant to our work. They achieve high-fidelity dynamic surface reconstruction from monocular RGB-D inputs. However, NDR falls short of realistic reconstruction in unobserved regions.
@@
Neural Rendering with Diffusion Priors. Recent years have witnessed remarkable progress in diffusion-based foundation models [50, 52], unlocking unprecedented capabilities for diverse applications such as image synthesis, editing, and shape generation [46]. DreamFusion [47] is a notable example that leverages pre-trained image diffusion priors to optimize a NeRF model through SDS, achieving high-quality text-to-3D generation. Building upon DreamFusion’s success, numerous follow-up works have emerged, particularly in text-to-3D [32, 12, 38, 69, 74, 81], image-to-3D [37, 77, 49, 64], scene editing [22, 53, 43], etc. Some other works [34, 56, 35] try to enhance the generation quality via fine-tuning diffusion prior. A noteworthy example is Zero-1-to-3 [34], which fine-tunes the stable-diffusion [50] model with relative camera pose conditioning. The resulting model takes an input image and the camera parameters of the viewpoint, and outputs an image rendered at the novel view. Recently, several works have been proposed for dynamic scene generation [58] and editing [53, 43]. MAV3D [58] adopts a text-to-video diffusion model [57] for text-to-4d generation. Control4D [53] utilizes a 2D diffusion prior to train a 4D GAN, enabling 4D editing of human portraits. AvatarStudio [43] fine-tunes the diffusion prior with time stamps and viewpoints, enabling editing of human head avatars with a view-and-time aware SDS loss. Different from these works, our focus is learning metrically accurate motion and geometry from the input monocular RGB-D video observations while achieving realistic completion in the unobserved region with a diffusion prior.
@@ 3.1 Dynamic Neural Surface Rendering
$\begin{gathered}C=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c}_{i},\quad D=\sum_{i=1}^{N}T_{i}\alpha_{i}d_{i},\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),\end{gathered}$ (4)
where $\alpha_{i}=1-\exp(-\sigma_{i}\delta_{i})$, $\delta_{i}$ is the distance between neighboring sample points.
@@
where ${\mathbf{I}}_{v}$ is the rendered image from the sampled virtual view, which has a relative pose $\Delta{\mathbf{P}}$ with respect to the real image observation ${\mathbf{I}}_{r}$ at the reference view. Here the relative pose $\Delta{\mathbf{P}}=[\Delta r,\Delta\omega,\Delta\phi]=[r_{v}-r_{r},\omega_{v}-\omega_{r},\varphi_{v}-\varphi_{r}]$ is parameterised in polar-coordinate. Conditioned on the reference view ${\mathbf{I}}_{r}$, the relative pose $\Delta{\mathbf{P}}$, and the sampled time-step $i\sim\mathcal{U}(0,1)$, the de-noising U-Net $\epsilon_{\psi}(\cdot)$ applies a de-noising step on the noisy latent $\tilde{{\mathbf{z}}}_{i}$ of the rendered view ${\mathbf{I}}_{v}$. $\theta$ is the learn-able parameters of our dynamic scene model. $w(i)$ is a time-dependent weighting term.
@@
The latent vectors of each keyframe for conditioning are pre-computed to save the computational cost. In practice, we find that using SDS does not have strict consistency across different frames, which affects the convergence of the deformation field. Thus, we incorporate a view-dependant modulation term to the weighting term $\hat{w}(i)=w(i)(\exp(|\Delta{\mathbf{P}}^{t}_{(1:2)}|)-1)$ to encourage the SDS loss to focus on novel views with larger view difference with respect to the reference view:
@@ 3.3 Optimization
Real View Supervision. Given a sampled input frame $I_{t}$, we randomly sample a batch of pixels and perform volume rendering by casting rays through those pixels and transforming the ray points to the world coordinate with the corresponding camera pose. The real-view supervision is formulated as:
@@
where $S$ is obtained by uniformly sampling $n$ points from $\hat{D}-tr/2$ to $\hat{D}+tr/2$ along sampled rays. The perturbed point ${\mathbf{x}}+\delta{\mathbf{x}}$ is sampled on a circle centered at ${\mathbf{x}}$ with radius $r$ orthogonal to the gradient vector $\nabla s({\mathbf{x}})$. Following [30] we compute the SDF gradient via finite difference.
Canonical Space Regularization. We find that performing regularization using all points on sampled rays in observation space leads to instability in learning the deformation of novel views. Instead, we propose to apply constraints directly to each individual slice of the hyper-dimensional canonical space by shortcutting the deformation network:
@@ 4.1 Experimental Setup
Method Metric KillingFusion [59] DeepDeform [8] iPhone [18]
Avg.
@@
NDR Acc. [cm] $\downarrow$
1.25
1.34
1.32
0.74
0.83
0.83
6.12
2.19
2.73
1.92
Comp. [cm] $\downarrow$
0.99
1.15
1.05
0.63
0.68
0.65
1.91
1.06
1.73
1.09
Clip sim. $\uparrow$
82.86
82.52
80.66
82.22
83.15
79.24
75.83
74.12
77.19
79.75
@@
Method Metric AMA [67] BANMo [78]
Avg.
Samba
Swing
eagle1
eagle2
NDR Acc. [cm] $\downarrow$
2.32
2.70
6.51
15.42
6.73
Comp. [cm] $\downarrow$
1.94
2.17
2.58
3.70
2.59
Clip sim. $\uparrow$
86.72
88.03
86.23
81.64
85.65
@@
Datasets. We evaluate our method on 9 real-world scenes that are from KillingFusion [59], DeepDeform [8], and iPhone dataset [18]. All videos in those datasets are captured with a consumer-grade RGB-D camera. Object masks are obtained with off-the-shelf object segmentation methods [13, 33] following NDR [9]. Additionally, we also evaluate our methods on 3 synthetic scenes from AMA [67] and BANMo [78] dataset which ground-truth meshes at all time-stamps are available. For AMA dataset [67], we render depth images from the calibrated 8 views aligned with the provided GT RGB images and masks, and select the front-view observations for optimisation. For BANMo dataset [78], we follow a similar protocol except that all the RGB, depth and mask are rendered. Following NDR [9], we also compensate for the object’s rigid motion and relative transformation in the data pre-processing stage. Please refer to our supplementary for more details.
Metrics. We adopt accuracy (acc. [cm]) and completion (comp. [cm]) to evaluate the surface reconstruction quality and CLIP similarity to evaluate the realism of completion. For 9 real-world scenes, we perform mesh culling to evaluate the observed region due to the lack of complete 360° GT meshes. The CLIP similarity is computed between RGB frames and the novel view renderings generated by our optimized model. These renderings follow a 360° trajectory around the object, which is in motion over time. In the case of the 3 synthetic scenes, we evaluate the complete mesh and render RGB images from the other 7 views to obtain CLIP similarity.
Implementation Details. We run MorpheuS on a desktop PC with an Intel Core i7-13700K CPU and NVIDIA RTX 4090 GPU with 24GB memory. The model takes around 2-3 hours to train and requires 10-22GB of memory depending on the the size of the input frames. We employ a coarse-to-fine training strategy that progressively increases the bandwidth of the positional encoding and hash encoding by a modulation term [45, 9, 73]. Additionally, we discover that when the learning rate is small, the learning of the canonical field (with hash encoding) is significantly faster than the deformation field (with positional encoding). This can let the model learn a good initialization of the canonical shape. Thus, we apply a learning rate warm-up strategy. In the warm-up phase, we freeze the deformation field when training the novel view with ${\mathcal{L}}_{\mathrm{S}}$ and set the timestep range of the diffusion prior to be $[0.02,0.5]$ to ensure the model learns a good canonical shape representation. In the second phase, we reduce the timestep range to $[0.02,0.2]$ to allow the model to learn accurate motion. The rendered resolution is doubled to improve the quality of the geometry and appearance completion. Please refer to supplementary material for more details.
@@
Table 3: Ablation studies on real-world datasets: a) w/o diffusion prior: optimize on real-view only; b) w/o depth: optimize with rgb only; c) w/o temporal condition: condition on first frame d) w/o angle weight: using original $w(i)$. e) w/o canonical space: modeling the scene in observation space via a quad-cube, which decomposes the 4D scene into 4 Hash feature grids.
@@ 4.2 Evaluation
Tab. 1 shows the per-scene quantitative results across three real-world datasets. For a more comprehensive evaluation, we extend our evaluations to include 4 synthetic scenes in Tab. 2. Both results show that our model demonstrates superior accuracy and completion in surface reconstruction, outperforming NDR which exhibits spurious surface completion due to the continuity MLPs (See Teddy in Fig. 3). Additionally, our method achieves a significantly better CLIP similarity in comparison to NDR, thanks to the effective knowledge distillation from the diffusion prior.
Fig. 3 shows a qualitative comparison of dynamic 360° reconstruction, highlighting the ability of diffusion priors to eliminate spurious surface completion and achieve photo-realistic reconstruction in unobserved regions. Notably, our model exhibits robustness in completing unobserved regions even in the presence of significant motion changes, as illustrated in the last row of the Teddy sequence.

### v2

@@ 1 Introduction
The challenge of reconstructing the dense 3D geometry, appearance, and motion of dynamic scenes from videos has persisted for decades, offering a wide range of applications in virtual reality and augmented reality. Traditional dense non-rigid reconstruction methods often assume strong priors on the object’s shape and motion due to the inherent ambiguities of the problem [89, 48]. Model-based approaches [87, 1, 28, 11, 76, 27] require 3D parametric models [2, 40, 49, 6, 31, 100] of specific object categories such as humans [2, 40], faces [6, 31], or animals [100], resulting in promising but category-constrained reconstructions. In this paper, we focus on model-agnostic approaches that could generalize to generic non-rigid objects. This task is often approached by jointly fusing a canonical shape and a per-frame deformation field from an RGB-D video sequence [47, 25, 69, 70, 8, 7]. Due to the constraints of physical sensors and the reliance on traditional voxel-grid representations, they often struggle to produce high-fidelity surface details and can only reconstruct observed regions.
Recent attention has shifted towards leveraging neural representations for scene reconstruction. Neural Radiance Fields (NeRF) [44] represent scene density and color in the weights of a neural network. In combination with volume rendering, NeRF achieves unprecedented performance on novel view synthesis. Many follow-up works adapt this idea into surface reconstruction [93, 84, 4, 83], SLAM [72, 98, 82, 29], dynamic scene reconstruction [56, 52, 53, 9] and more. Notably, NDR [9] adopts NeRF-style representations for dynamic RGB-D surface reconstruction. Thanks to the expressiveness of neural representations, they achieve accurate and smooth surface reconstruction. However, real-world captures often contain large unobserved regions. The inherent smoothness of neural representations, and the difficulty to incorporate domain-specific priors, cause current methods to fall short of achieving realistic completion.
@@
Our key contribution is to demonstrate the capability to learn metrically accurate geometry and deformations of dynamic objects from casually captured RGB-D videos while achieving realistic completion in unobserved regions with diffusion priors. We propose a temporal view-dependent SDS strategy to improve the realism of completion while learning an accurate deformation field. A canonical space regularization strategy is used to avoid the trivial solution of surface completion. To the best of our knowledge, MorpheuS is the first to achieve accurate, photo-realistic 360°$\degree$ surface reconstruction of an arbitrary dynamic object from casually captured monocular RGB-D video.
@@ 2 Related Work
Dense Non-rigid Reconstruction. Traditional dense dynamic scene reconstruction methods often require strong prior knowledge about the object shape and motion, fitting 3D prior parametric models [40, 61, 100, 31, 87, 1, 28, 11, 76, 27] to video sequences of a specific object category, and thus cannot work with arbitrary unseen objects. Early model-agnostic methods either require accurate dense correspondences [20] or a template shape of the object to be reconstructed first before tracking the dynamic motion [73, 99, 95, 5]. The seminal work DynamicFusion [47] was the first direct and template-free method to achieve real-time simultaneous tracking and reconstruction of generic non-rigid objects via joint optimization of a canonical shape and a per-frame deformation field. Follow-up methods improve the robustness of motion tracking and reconstruction quality by incorporating photo-metric constraints [21], classical  [25] or learned correspondences [8, 7], reformulating the optimization as direct dense SDF alignment [69, 70]. Although these methods have shown promising reconstruction results, they fail to complete the unobserved regions due to their inherent traditional scene representation.
@@
Neural Rendering for Dynamic Scenes. The success of NeRF [44] in novel view synthesis has inspired many follow-up works [56, 52, 32, 53, 18, 77, 79, 16, 9, 19, 91, 92, 34, 71, 45, 43] in dynamic scene reconstruction. These approaches can be broadly classified into two paradigms: a) Directly modeling scene geometry and motion as a unified 4D space-time field in world (observation) space [32, 18, 10, 17, 34, 64, 24], where motion priors, such as scene flow or trajectory can be involved as an additional regularization [32, 79, 34]. b) Learning a canonical shape by warping the shape from observation space to canonical space via a separate deformation field [56, 52, 53, 77, 16, 9, 19, 88, 91, 92, 71, 45, 15]. Among these works, NDR [9] is particularly relevant to our work. They achieve high-fidelity dynamic surface reconstruction with neural representations. However, the continuity of neural representations does not allow realistic completion in unobserved regions.
@@
Neural Rendering with Diffusion Priors. Recent years have witnessed remarkable progress in diffusion-based foundation models [60, 62], unlocking unprecedented capabilities for diverse applications such as image synthesis, editing, and shape generation [54]. DreamFusion [55] is a notable example that leverages pre-trained image diffusion priors to optimize a NeRF model through SDS, achieving high-quality text-to-3D generation. Building upon DreamFusion’s success, numerous follow-up works have emerged, particularly in text-to-3D [35, 12, 42, 81, 86, 97], image-to-3D [41, 90, 57, 75], scene editing [23, 63, 50], etc. Some other works [38, 66, 39, 30] try to improve the generation quality via fine-tuning pre-trained diffusion models. A noteworthy example is Zero-1-to-3 [38], which fine-tunes stable-diffusion [60] with relative pose conditioning.
Concurrent works. Recently, several works have been proposed for dynamic scene generation [68, 26, 59, 37, 94, 51, 80] and editing [63, 50]. MAV3D [68] adopts a text-to-video diffusion model [67] for text-to-3D generation. Control4D [63] utilizes a 2D diffusion prior to train a 4D GAN for 4D editing of human portraits. AvatarStudio [50] fine-tunes the diffusion prior with time stamps and viewpoints, enabling editing of human head avatars with a view-and-time aware SDS loss. Consistent4d [26], Efficient4D [51], and DpDy [80] use diffusion prior for RGB-based dynamic scene reconstruction. Different from these works, our focus is learning metrically accurate motion and geometry from monocular RGB-D video while achieving realistic completion in the unobserved region with a diffusion prior.
@@ 3.1 Dynamic Neural Surface Rendering
$\begin{gathered}C=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c}_{i},\quad D=\sum_{i=1}^{N}T_{i}\alpha_{i}d_{i},\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right).\end{gathered}$ (4)
Here $T_{i}$ represents transmittance. The alpha value $\alpha_{i}$ is calculated as $1-\exp(-\sigma_{i}\delta_{i})$, where $\delta_{i}$ denotes the distance between neighboring sample points.
@@
where ${\mathbf{I}}_{v}$ is the rendered image from the sampled virtual view, which has a relative pose $\Delta{\mathbf{P}}$ with respect to the real image observation ${\mathbf{I}}_{r}$ at the reference view. Here the relative pose $\Delta{\mathbf{P}}=[\Delta r,\Delta\omega,\Delta\phi]=[r_{v}-r_{r},\omega_{v}-\omega_{r},\varphi_{v}-\varphi_{r}]$ is parameterized in polar-coordinate. Conditioned on the reference view ${\mathbf{I}}_{r}$, the relative pose $\Delta{\mathbf{P}}$, and the sampled time-step $i\sim\mathcal{U}(0,1)$, the de-noising U-Net $\epsilon_{\psi}(\cdot)$ applies a de-noising step on the noisy latent $\tilde{{\mathbf{z}}}_{i}$ of the rendered view ${\mathbf{I}}_{v}$. $\theta$ is the learn-able parameters of our dynamic scene model. $w(i)$ is a time-dependent weighting term.
@@
The latent vectors of each keyframe for conditioning are pre-computed to save the computational cost. To mitigate the negative effects of view-inconsistency stemming from the SDS loss, inspired by [74], we incorporate a view-dependant modulation term to the weighting term $\hat{w}(i)=w(i)(\exp(|\Delta{\mathbf{P}}^{t}_{(1:2)}|)-1)$. This reduces the gradient for virtual views that have larger view differences with respect to the reference view:
@@ 3.3 Optimization
Real View Supervision. Given a sampled input frame $I_{t}$, we randomly sample a batch of pixels and cast rays through those pixels to perform volume rendering. The real-view supervision is formulated as:
@@
where $S$ is obtained by uniformly sampling $n$ points from $\hat{D}-tr/2$ to $\hat{D}+tr/2$ along sampled rays. The perturbed point ${\mathbf{x}}+\delta{\mathbf{x}}$ is sampled on a circle centered at ${\mathbf{x}}$ with radius $r$ orthogonal to the gradient vector $\nabla s({\mathbf{x}})$. Following [33], we compute the SDF gradient via finite difference.
Canonical Space Regularization. We find that performing regularization using all points on sampled rays in observation space may lead to degenerate solutions. Instead, we propose to apply constraints directly to each individual slice of the hyper-dimensional canonical space by skipping the deformation network:
@@ 4.1 Experimental Setup
Method Metric KillingFusion [69] DeepDeform [8] iPhone [19] Avg.
@@
NDR Acc. [cm] $\downarrow$ 1.25 1.34 1.32 0.74 0.83 0.83 6.12 2.19 2.73 1.92
Comp. [cm] $\downarrow$ 0.99 1.15 1.05 0.63 0.68 0.65 1.91 1.06 1.73 1.09
Clip sim. $\uparrow$ 82.86 82.52 80.66 82.22 83.15 79.24 75.83 74.12 77.19 79.75
@@
Method Metric AMA [78] BANMo [91] Avg.
Samba Swing Eagle1 Eagle2
NDR Acc. [cm] $\downarrow$ 2.32 2.70 6.51 15.42 6.73
Comp. [cm] $\downarrow$ 1.94 2.17 2.58 3.70 2.59
Clip sim. $\uparrow$ 86.72 88.03 86.23 81.64 85.65
@@
Datasets. We evaluate our method on 9 real-world scenes that are from KillingFusion [69], DeepDeform [8], and iPhone dataset [19]. All videos in those datasets are captured with a consumer-grade RGB-D camera. Object masks are obtained with off-the-shelf object segmentation methods [13, 36, 14]. Additionally, we also evaluate our methods on 3 synthetic scenes from AMA [78] and BANMo [91] dataset which ground-truth meshes at all time-stamps are available. For AMA dataset [78], we render depth images from the calibrated 8 views aligned with the provided GT RGB images and masks and select the front-view observations for optimisation. For BANMo dataset [91], we follow a similar protocol except that all the RGBs, depths and masks are rendered. Following NDR [9], we compensate for the object’s rigid motion and relative transformation in the data pre-processing stage using Robust-ICP [96].
Metrics. We adopt accuracy (acc. [cm]) and completion (comp. [cm]) to evaluate the surface reconstruction quality and CLIP similarity [58] to evaluate the realism of completion. For 9 real-world scenes, we perform mesh culling for evaluation due to the lack of complete GT meshes. The CLIP similarity is computed between RGB frames and the novel view renderings generated by our optimized model. These renderings follow a 360° trajectory around the object, which is in motion over time. For 3 synthetic scenes, we evaluate the complete mesh and render RGB images from the other 7 views to obtain CLIP similarity.
Implementation Details. We run MorpheuS on a desktop PC with an Intel Core i7-13700K CPU and NVIDIA RTX 4090 GPU with 24GB memory. The model takes around 2-3 hours to train and requires 10-22GB of memory depending on the the size of the input frames. We employ a coarse-to-fine training strategy that progressively increases the number of frequency bands of the positional encoding and hash encoding by a modulation term [53, 9, 85]. Additionally, we discover that when the learning rate is small, the learning of the canonical field (with hash encoding) is significantly faster than the deformation field (with positional encoding). This can let the model learn a good initialization of the canonical shape. Thus, we apply a learning rate warm-up strategy. In the warm-up phase, we freeze the deformation field when training the novel view with ${\mathcal{L}}_{\mathrm{S}}$ and set the timestep range of the diffusion prior to be $[0.02,0.5]$ to ensure the model learns a good canonical shape representation. In the second phase, we reduce the timestep range to $[0.02,0.2]$ to allow the model to learn accurate motion. The rendered resolution is doubled to improve the quality of the geometry and appearance completion.
@@
Table 3: Ablation studies on real-world datasets: a) w/o diffusion prior: optimize on real-view only; b) w/o depth: optimize with rgb only; c) w/o temporal condition: condition on first frame d) w/o angle weight: using original $w(i)$. e) w/o canonical space: model the scene as a unified 4D space-time field in world space.
@@ 4.2 Evaluation
Tab. 1-2 show the per-scene quantitative results on 9 real-world scenes and 4 synthetic scenes respectively. Both results show that our model demonstrates superior accuracy and completion in surface reconstruction, outperforming NDR which exhibits spurious surface completion due to the continuity MLPs (See Teddy in Fig. 3). Additionally, our method achieves a significantly better CLIP similarity in comparison to NDR, thanks to the effective knowledge distillation from the diffusion prior.
Fig. 3 shows a qualitative comparison of dynamic 360° reconstruction, highlighting the ability of diffusion priors to eliminate spurious surface prediction and achieve photo-realistic completion. Notably, our model exhibits robustness in completing unobserved regions even in the presence of significant motion changes, as illustrated in the last row of the Teddy sequence. Fig 4 shows a qualitative result of novel view synthesis. Training dynamic neural representations within a monocular setting typically leads to poor quality novel view synthesis in the case of large viewpoint differences. In contrast, our diffusion prior helps to alleviate this issue allowing 360° novel view synthesis. Please refer to supplementary materials for further quantitative analysis on novel view synthesis.